40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  ATF eLog  Not logged in ELOG logo
Entry  Fri Dec 13 01:51:15 2019, Raymond, DailyProgress, PD QE, Cryostat wiring fixes 18_32malepinsinairAssigned.pdf
    Reply  Mon Dec 16 18:16:31 2019, Raymond, DailyProgress, PD QE, Cryostat wiring fixes 
Message ID: 2480     Entry time: Mon Dec 16 18:16:31 2019     In reply to: 2479
Author: Raymond 
Type: DailyProgress 
Category: PD QE 
Subject: Cryostat wiring fixes 

Opened the cryostat to resolder the heater and re-wrap the thermal anchor for the sample RTD and PD connection. All connections are working as expected at room temperature. A2 is still in the sample mount. 

ELOG V3.1.3-